The displacement x variation of a particle with time t is given by expression x is equal to - 15t square + 20t + 30 find the position of the body at t is equal to zero second and t is equal to 6 seconds if this particle moving with uniform acceleration ?explain.
1
Expert's answer
2018-04-24T08:58:08-0400
Find the position of the particle at t = 0 s:
x(0)=-15×0+20×0+30=30 m

Find the position of the particle at t = 6 s:
x(6)=-15×6^2+20×6+30=-390 m